What volume of a 15% by mass NaOH, which its density is 1.116g/mL, should be used to make 5.40L of an NaOH solution with a pH of 10.2?
Answer #1

pH = 10.2
pOH = 14 - pH = 14 - 10.2 = 3.8

[OH-] = 1 x 10-pH = 1 x 10-3.8 = 1.58 x 10-4 M

[NaOH] = [OH-] = 1.58 x 10-4 M

moles of NaOH = molarity of NaOH x volume of NaOH = 1.58 x 10-4 x 5.40 = 8.56 x 10-4 mol

mass of NaOH solute = moles of NaOH x molar mass of NaOH = 8.56 x 10-4 x 40 = 0.034 g

mass of 15% NaOH solution = mass of NaOH salute / 15% = 0.034 / 0.15 = 0.227 g

volume of 15% NaOH = mass of 15% NaOH solution / density of 15% NaOH solution

= 0.227 g/ 1.116 g/ml = 0.203 ml